OpenStudy (anonymous):

PLEASE HELP ! Find all critical points of the following functions and determine whether there is a local maximum, a local minimum, or a saddle point at those points. f(x;y)=(x+y)(xy+1) [Hint: there are two critical points]

OpenStudy (anonymous):

Do you know how to take paritial derivatives?

OpenStudy (anonymous):

I would probably multiiply everything out, then take the first and second partial derivatives of the function with respect to x, then take the first and second partial derivatives of the function with respect to y, and then take the partial derivative WRT x then WRT y.

OpenStudy (anonymous):

@calmat01 yes what i did was Fx= 2xy +1 y'2 Fy= x^2 2yx + 1 but when you get the critical points... how!?

OpenStudy (anonymous):

Whoa, a partial with respect to x means the y terms are constant. So, your \[f _{x}(x,y)=2xy+1+y ^{2}\] and then the partial WRT y is going to be \[f _{y}(x,y)=x ^{2}+2xy+1\]

OpenStudy (anonymous):

So, your partial WRT y is correct. The partial WRT x needs to be amended.

OpenStudy (anonymous):

oops typo Fx= 2xy +1 + y^2 Fy= x^2 + 2yx + 1

OpenStudy (anonymous):

so how do i find the critical numbers?

OpenStudy (anonymous):

Now, you set each partial derivative = 0, and solve. Since we need to find the points that will satisfy both simultaneously, set them equal to each other and the 1 + 2xy terms will cancel out, leaving you with \[x ^{2}=y ^{2}\]

OpenStudy (anonymous):

so it's (0,0) ?

OpenStudy (anonymous):

@calmat01

OpenStudy (anonymous):

You will find that (0,0) will not make either partial derivative 0.

OpenStudy (anonymous):

Try something like (-1, 1) or (1,-1)

OpenStudy (anonymous):

howw!? :((

OpenStudy (anonymous):

We know that x^2=y^2, based on setting our partial derivatives =0, right?

OpenStudy (anonymous):

So, we know that \[x ^{2}-y ^{2}=0. \] But this factors as the difference of two squares. It factors as (x-y)(x+y)=0, which tells us that either x=y, or x=-y. Now go back into your partial derivatives and replace x with y or replace x with -y and you will find the values I came up with. Try it!

OpenStudy (anonymous):

@alyannahere did you get it?

OpenStudy (anonymous):

Those are not your only possibilities, but I was just giving you the two that worked to make your partials=0.

OpenStudy (anonymous):

righttt. i think i got it? hahahah

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

Most welcome.